OpenStudy (anonymous):

A mass is moving in the positive x direction with constant acceleration a(t)=-9.8. Find v(t) if v(0)=3 and Find s(t) if v(0)=3 and s(0)=4. For v(t) I got -4.9t^(2) + C. But for the second part, I am confused.

OpenStudy (anonymous):

these are the notes I found. v'(t)=a(t), v(0)=initial velocity and s'(t)=v(t), s(0)= initial position

OpenStudy (xapproachesinfinity):

yes we have a=v' so we need to integrate a(t)=-9.8

OpenStudy (xapproachesinfinity):

how did you find v(t)=-4.9t+c?

OpenStudy (xapproachesinfinity):

it is correct but do you know the method

OpenStudy (xapproachesinfinity):

to find the constant c we know that v(0)=3 so v(0)=c=3 so v(t)=-4.9t+3 the method you used to get v is the same one to get s(t) where v(t)=s'(t)

OpenStudy (xapproachesinfinity):

oh hold on that's wrong v(t)=-9.8t+c not t^2 so c=3 v(t)=-9.8t+3

OpenStudy (xapproachesinfinity):

t^2 will be in the position function not v it will be s(t)

OpenStudy (anonymous):

oh okay thank you for correcting my first part

OpenStudy (anonymous):

as for the second, let me see...I think i still need help though

OpenStudy (anonymous):

basically s'(t) will be a constant, yeah?

OpenStudy (anonymous):

correction in instruction s(0)=4

OpenStudy (anonymous):

is s(t) then just equal to 4?